Ray is 14 years older than his sister in 3 years he will be twice as old how old is each of them
Answer by stanbon(75887)

Ray is 14 years older than his sister. In 3 years he will be twice as old.
How old is each of them?
====
r = s + 14
r+3 = 2(s+3)
-----
Substitute for "r" and solve for "s"::
s+14+3 = 2s+6
s = 11 yrs (sister's age now)
------
Solve for "r"::
r = s + 14
r = 11+14 = 25 yrs (Ray's age now)
-----------
Cheers,
Stan H.
---------